OpenStudy (anonymous):

Write an equation in slope-intercept form of a line that contains the point (6,-5) and is perpendicular to the equation of the line y=-3x+5 *Hint*(you may use the point-slope form to get started)

OpenStudy (anonymous):

first find the slope perpendicular to the equation of that line. remember that perpendicular slopes multiply to be -1 so for instance in your case (-3)x=-1 where x will be the slope perpendicular. next you need to finish the equation by putting your point into point slope form. this is y-(y of point)=(your perpendicular slope)*(x-(x of point)) that should get you the answer you need after you simplify

OpenStudy (anonymous):

wouldn't the perpendicular equation just be y=1/3x+5?

OpenStudy (anonymous):

when i put it into point-slope i got y=1/3x-7.. not sure if that was right though?

OpenStudy (anonymous):

your answer should end up something along the lines of y=(1/3)x+(23/3) if i did my math right

OpenStudy (anonymous):

23/3? how..

OpenStudy (anonymous):

y-(-5)=1/3(X-6) y+5=1/3x-2 -5 5 y=1/3x-7

OpenStudy (anonymous):

haha yeah i did my math wrong. long night you should be right

OpenStudy (anonymous):

lol :p so its y=1/3x-7?

OpenStudy (anonymous):

that looks right to me

OpenStudy (anonymous):

cool, thanks c:

OpenStudy (anonymous):

welcome